OpenStudy (anonymous):

Help Plz will medal and fan :) What is the slope-intercept form of the function that contains the point (6, 2) and has a slope of 3? y=___x+___ @Luigi0210 @Godlovesme

OpenStudy (anonymous):

use the formula y=mx+b

OpenStudy (anonymous):

then substitute

OpenStudy (godlovesme):

we know that slope is 3 right? @FateRocks8D

OpenStudy (anonymous):

yes

OpenStudy (godlovesme):

y=mx+b m represents the slope so, plug that

OpenStudy (anonymous):

y=3x+b ?

OpenStudy (godlovesme):

yep :)

OpenStudy (godlovesme):

now we need b which is y-intercept -> is where the line hits the y-axis so basically x will be 0 during y-intercepts

OpenStudy (godlovesme):

r u confused? @FateRocks8D :P

OpenStudy (anonymous):

lil bit

OpenStudy (godlovesme):

sorry :( which part is confusing u?

OpenStudy (anonymous):

for b

OpenStudy (godlovesme):

to make this easier use this and solve for b 2=3(6) + b 2-12=12-12+b b=-10 y=3x-10 is that an option?

OpenStudy (anonymous):

so u pluged in 6,2 and added and suib. and got -10

OpenStudy (godlovesme):

yep :)

OpenStudy (anonymous):

ok cuz the last time I had a question like this I failed it o_o

OpenStudy (godlovesme):

:(